OpenStudy (anonymous):

The radioactive element americium-241 has a half-life of 432 years. Suppose we start with a 20-g mass of americium-241. How much will be left after 349 years? Compute the answer to three significant digits.

OpenStudy (anonymous):

1/2= e^-(c*432) c= ?

OpenStudy (unklerhaukus):

\[[\text{Am}]_t=[\text{Am}]_02^{{-t}/{t_{½}}}\]\[[\text{Am}]_0=20[\text{g}]\qquad t_{1/2}=432[\text{yr}]\]\[[\text{Am}]_t=20[\text{g}]2^{{-t}/{432[\text{yr}]}}\]\[t=349[\text{yr}]\]\[[\text{Am}]_t=20[\text{g}]2^{{-349[\text{yr}]}/{432}[\text{yr}]}\]